# Generalized Zykov’s Theorem

### v1

@@ Abstract
For a simple graph $G$, let $n$ denote its number of vertices, and let $N(G,K_{t})$ denote the number of copies of $K_{t}$ in $G$. Zykov’s theorem asserts that for any $K_{r+1}$-free graph and $t\geq 2$,
@@ 1.1 Localization
We use a similar argument to give a vertex-based localization of 1.2.
@@ 2.1.2 Characterizing Extremal Graphs
Thus assume $G[Supp(z)]$ is not a clique. Therefore, there exists $i,j\in Supp(z)$ such that $i$ and $j$ are not adjacent. From 1, we have $\Phi(G,y)=\Phi(G,z)$, where $y=T_{z_{j}}^{ij}(z)$. Note that $Supp(y)=Supp(z)\setminus\{j\}$, therefore $Supp(y)\subsetneq Supp(z)$. Since, $y$ is a minimizer of $\Phi(G,\cdot)$, from our assumption we get that $G[Supp(y)]$ is a complete $\omega_{y}$-partite graph.
Let $P_{1},P_{2},\dots,P_{\omega_{y}}$ be the parts in $G[Supp(y)]$.
@@
Assume on the contrary, there exists $v_{k}\in N(j)\cap P_{k}$ for all $k\in[\omega_{y}]$. Note that we can define $y^{*}=(y^{*}_{1},y^{*}_{2},\dots,y^{*}_{n})\in\Delta^{n-1}$ such that $y^{*}_{v_{k}}=\sum_{v\in P_{k}}y_{v}$ for all $k\in[\omega_{y}]$ and $y^{*}_{l}=0$ for all $l\notin\{v_{1},v_{2},\dots,v_{\omega_{y}}\}$ (that is, for $k\in[\omega_{y}]$, we transfer all the amount from the coordinates indexed by $P_{k}$ to the coordinate corresponding $v_{k}$). Since each $P_{k}$ is an independent set, from 2, we get that $\Phi(G,y^{*})=\Phi(G,y)$. Clearly, $Supp(y^{*})$ induces a $\omega_{y}$-order clique and $y^{*}$ is a minimizer. Thus, from 2.2 we get that $y^{*}$ is a minimal support minimizer. Therefore, using equality condition 3 we have $c(v_{k})=|Supp(y^{*})|=\omega_{y}$. From the assumption we get that $Supp(y^{*})\cup\{j\}$ induces an $\omega_{y}+1$-order clique. Since $v_{k}$ is contained in this clique, we have $c(v_{k})\geq\omega_{y}+1$, which results a contradiction. ∎
@@
Suppose there exists $k\neq 1$, such that $P_{k}\cap N(j)=\emptyset$.
First suppose $|P_{1}|>1$. Let $y^{\prime}=T_{z_{i}}^{ji}(z)$. Clearly from 1, $\Phi(G,y^{\prime})=\Phi(G,z)$ and $|Supp(y^{\prime})|=|Supp(z)|-1$. Thus, from induction hypothesis. $G[Supp(y^{\prime})|$ is a complete $\omega_{y^{\prime}}$-partite graph. Clearly $P^{\prime}_{1}=P_{1}\setminus\{i\},P_{2},\dots,P^{\prime}_{k}=P_{k}\cup\{j\},\dots,P_{\omega_{y}}$ are maximal independent sets in $G[Supp(y^{\prime})]$. It is easy to see that $\{P^{\prime}_{1},P_{2},\dots,P^{\prime}_{k},\dots,P_{\omega_{y}}\}$ form the parts in $G[Supp(y^{\prime})]$, therefore $\omega_{y^{\prime}}=\omega_{y}$. Note that $y^{\prime}_{v}=y_{v}$ for all $v\notin\{i,j\}$, $y^{\prime}_{i}=y_{j}=0$, and $y^{\prime}_{j}=y_{i}=z_{i}+z_{j}$. From the induction hypothesis we have $\sum_{v\in P_{1}}y_{v}=\sum_{v\in P_{k}}y_{v}=\frac{1}{\omega_{y}}$. But then we get $\sum_{v\in P^{\prime}_{1}}y^{\prime}_{v}<\sum_{v\in P^{\prime}_{k}}y^{\prime}_{v}$, which results in a contradiction.
Now suppose $|P_{1}|=1$. Clearly $\{P_{2},\dots,P^{\prime}_{k},\dots,P_{\omega_{y}}\}$ are the parts in $G[Supp(y^{\prime})]$. Suppose that $G[\operatorname{Supp}(y^{\prime})]$ is $1$-partite, that is, $P_{k}^{\prime}$ is the only part. By transferring the weight of all vertices in $P_{k}^{\prime}$ to a single vertex $v\in P_{k}^{\prime}$, 2 implies that we obtain another minimizer whose support consists solely of $v$. Then, by equality condition 3, we have $c(v)=1$. However, this contradicts our assumption that $G$ has no isolated vertices.
Thus, $P^{\prime}_{k}$ is not the only part in $G[Supp(y^{\prime})]$. From the induction hypothesis we have $\sum_{v\in P_{t}}y_{v}=\sum_{v\in P_{k}}y_{v}$, where $t\notin\{1,k\}$. But then we get $\sum_{v\in P_{t}}y^{\prime}_{v}<\sum_{v\in P^{\prime}_{k}}y^{\prime}_{v}$, which results in a contradiction. ∎
Thus, we get that $P_{k}\cap N(j)\neq\emptyset$ for all $k\neq 1$.
@@
Suppose for some $k\neq 1$, there exists $v_{k}\in P_{k}$ such that $v_{k}$ is not adjacent to $j$. But from 2, $P_{k}\cap N(j)\neq\emptyset$. Thus there exists $u_{k}\in P_{k}$ which is adjacent to $j$. In $G[Supp(y^{\prime})]$, where $y^{\prime}=T_{x_{i}}^{ji}(z)$, $v_{k}$ and $u_{k}$ are in the same part but $N(v_{k})\neq N(u_{k})$. This contradicts the fact that $G[Supp(y^{\prime})]$ is complete multipartite graph. ∎
Thus from 3 we get that $j$ is adjacent to all the vertices in $\bigcup_{i=2}^{\omega_{y}}P_{i}$. Also from 2 we get that $j$ is not adjacent to any vertex in $P_{1}$. Therefore, $G[Supp(z)]$ is a complete multipartite graph, with parts $V_{1}=P_{1}\cup\{j\},V_{2}=P_{2},\dots,V_{\omega_{y}}=P_{\omega_{y}}$. Note that $\omega_{z}=\omega_{y}$.
From the induction Hypothesis we know that, for all $k\in[\omega_{z}]$, $\sum_{v\in P_{k}}y_{v}=\frac{1}{\omega_{y}}$. We have, $y_{v}=z_{v}$ for all $v\notin\{i,j\}$, $y_{i}=z_{i}+z_{j}$, and $y_{j}=0$. Note that $i,j\in V_{1}$, therefore $\sum_{v\in V_{k}}y_{v}=\sum_{v\in V_{k}}z_{v}$ for all $k\in[\omega_{z}]$. Thus, for all $k\in[\omega_{z}]$,
$\sum_{v\in V_{k}}z_{v}=\frac{1}{\omega_{y}}=\frac{1}{\omega_{z}}$
From the equality condition 1, we obtain that $y=\left\{\frac{1}{n},\frac{1}{n},\dots,\frac{1}{n}\right\}$ is a minimizer of $\Phi(G,\cdot)$. Hence, by 2.3, the graph $G$ must be complete multipartite. Let $V_{p}$ and $V_{q}$ be any two parts of $G$. Using 2.3 again, we have

### v2

@@ Abstract
For a simple graph $G$, let $n$ denote its number of vertices, and let $N(G,K_{t})$ denote the number of copies of $K_{t}$ in $G$. Zykov’s theorem (1949) asserts that for any $K_{r+1}$-free graph and $t\geq 2$,
@@ 1.1 Localization
We adapt this argument to give a vertex-based localization of 1.2. The challenge here is to design the appropriate function that works in this setting.
@@ 2.1.2 Characterizing Extremal Graphs
Thus assume $G[Supp(z)]$ is not a clique. Therefore, there exists $i,j\in Supp(z)$ such that $i$ and $j$ are not adjacent. From 1, we have $\Phi(G,y)=\Phi(G,z)$, where $y=T_{z_{j}}^{ij}(z)$. Note that $Supp(y)=Supp(z)\setminus\{j\}$, therefore $Supp(y)\subsetneq Supp(z)$. Since, $y$ is a minimizer of $\Phi(G,\cdot)$, from our assumption we get that $y$ satisfies the lemma and therefore $G[Supp(y)]$ is a complete $\omega_{y}$-partite graph. Let $P_{1},P_{2},\dots,P_{\omega_{y}}$ be the parts in $G[Supp(y)]$.
@@
Assume on the contrary, there exists $v_{k}\in N(j)\cap P_{k}$ for all $k\in[\omega_{y}]$. Note that we can define $y^{*}=(y^{*}_{1},y^{*}_{2},\dots,y^{*}_{n})\in\Delta^{n-1}$ such that $y^{*}_{v_{k}}=\sum_{v\in P_{k}}y_{v}$ for all $k\in[\omega_{y}]$ and $y^{*}_{l}=0$ for all $l\notin\{v_{1},v_{2},\dots,v_{\omega_{y}}\}$ (that is, for $k\in[\omega_{y}]$, we transfer all the amount from the coordinates indexed by $P_{k}$ to the coordinate corresponding $v_{k}$). Therefore, $Supp(y^{*})=\{v_{1},v_{2},\dots v_{\omega_{y}}\}.$ Since each $P_{k}$ is an independent set, from 2, we get that $\Phi(G,y^{*})=\Phi(G,y)$. Thus, $y^{*}$ is a minimizer. Clearly, $Supp(y^{*})$ induces an $\omega_{y}$-order clique. Therefore, from 2.2 we get that $y^{*}$ is a minimal support minimizer. From equality condition 3 we have $c(v_{k})=|Supp(y^{*})|=\omega_{y}$. But from the assumption we get that $Supp(y^{*})\cup\{j\}$ induces an $\omega_{y}+1$-order clique. Since $v_{k}$ is contained in this clique, we have $c(v_{k})\geq\omega_{y}+1$, which results a contradiction. ∎
@@
Suppose there exists $k>1$, such that $P_{k}\cap N(j)=\emptyset$.
First suppose $|P_{1}|>1$. Let $y^{\prime}=T_{z_{i}}^{ji}(z)$. Clearly from 1, $\Phi(G,y^{\prime})=\Phi(G,z)$, that is, $y^{\prime}$ is also a minimizer. Moreover, $Supp(y^{\prime})\subsetneq Supp(z)$. Thus, from our assumption, the lemma holds for $y^{\prime}$. Therefore, $G[Supp(y^{\prime})|$ is a complete $\omega_{y^{\prime}}$-partite graph. Clearly $P^{\prime}_{1}=P_{1}\setminus\{i\},P_{2},\dots,P^{\prime}_{k}=P_{k}\cup\{j\},\dots,P_{\omega_{y}}$ are maximal independent sets in $G[Supp(y^{\prime})]$. It is easy to see that $\{P^{\prime}_{1},P_{2},\dots,P^{\prime}_{k},\dots,P_{\omega_{y}}\}$ form the parts in $G[Supp(y^{\prime})]$, therefore $\omega_{y^{\prime}}=\omega_{y}$. Note that $y^{\prime}_{v}=y_{v}$ for all $v\notin\{i,j\}$, $y^{\prime}_{i}=y_{j}=0$, and $y^{\prime}_{j}=y_{i}=z_{i}+z_{j}$. As $y$ satisfies the lemma, $\sum_{v\in P_{1}}y_{v}=\sum_{v\in P_{k}}y_{v}=\frac{1}{\omega_{y}}$. But then we get $\sum_{v\in P^{\prime}_{1}}y^{\prime}_{v}<\sum_{v\in P^{\prime}_{k}}y^{\prime}_{v}$, which results in a contradiction, since the lemma holds for $y^{\prime}$ and therefore, $\sum_{v\in P^{\prime}_{1}}y^{\prime}_{v}=\sum_{v\in P^{\prime}_{k}}y^{\prime}_{v}$.
Now suppose $|P_{1}|=1$. Clearly $\{P_{2},\dots,P^{\prime}_{k},\dots,P_{\omega_{y}}\}$ are the parts in $G[Supp(y^{\prime})]$, where $P^{\prime}_{k}=P_{k}\cup\{j\}$. Suppose that $G[\operatorname{Supp}(y^{\prime})]$ is $1$-partite, that is, $P_{k}^{\prime}$ is the only part. Using 2, by transferring the total amount from the coordinates indexed by $P_{k}^{\prime}$ to the coordinate corresponding to a single vertex $v\in P_{k}^{\prime}$, we obtain another minimizer whose support consists solely of $v$. Then, by equality condition 3, we have $c(v)=1$, therefore, $v$ is an isolated vertex. However, this contradicts our assumption that $G$ has no isolated vertices.
Thus, $P^{\prime}_{k}$ is not the only part in $G[Supp(y^{\prime})]$. Since the lemma holds for $y$, we have $\sum_{v\in P_{t}}y_{v}=\sum_{v\in P_{k}}y_{v}$, where $t\notin\{1,k\}$. Note that, $\sum_{v\in P_{t}}y_{v}=\sum_{v\in P_{t}}y^{\prime}_{v}$ and $\sum_{v\in P_{k}}y_{v}<\sum_{v\in P^{\prime}_{k}}y^{\prime}_{v}$. Thus we get, $\sum_{v\in P_{t}}y^{\prime}_{v}<\sum_{v\in P^{\prime}_{k}}y^{\prime}_{v}$, which results in a contradiction, since $y^{\prime}$ must also satisfy the lemma. ∎
From 1, there exists $k\in[\omega_{y}]$ such that $N(j)\cap P_{k}=\emptyset$, and from 2, for $k>1$, $N(j)\cap P_{k}\neq\emptyset$. Thus $N(j)\cap P_{1}=\emptyset$.
@@
Suppose for some $k>1$, there exists $v_{k}\in P_{k}$ such that $v_{k}$ is not adjacent to $j$. But from 2, $P_{k}\cap N(j)\neq\emptyset$. Thus there exists $u_{k}\in P_{k}$ which is adjacent to $j$. In $G[Supp(y^{\prime})]$, where $y^{\prime}=T_{x_{i}}^{ji}(z)$, $v_{k}$ and $u_{k}$ are in the same part but $N(v_{k})\neq N(u_{k})$. This contradicts the fact that $G[Supp(y^{\prime})]$ is a complete multipartite graph. ∎
Thus from 3 we get that $j$ is adjacent to all the vertices in $\bigcup_{i=2}^{\omega_{y}}P_{i}$. Also we know that $N(j)\cap P_{1}=\emptyset$. Therefore, $G[Supp(z)]$ is a complete multipartite graph, with parts $V_{1}=P_{1}\cup\{j\},V_{2}=P_{2},\dots,V_{\omega_{y}}=P_{\omega_{y}}$. Note that $\omega_{z}=\omega_{y}$.
Since $y$ satisfies the lemma, we know that, for all $k\in[\omega_{z}]$, $\sum_{v\in P_{k}}y_{v}=\frac{1}{\omega_{y}}$. Note that $i,j\in V_{1}$ and we have, $y_{v}=z_{v}$ for all $v\notin\{i,j\}$. Therefore, for all $k>1$ we get,
$\sum_{v\in V_{k}}z_{v}=\sum_{v\in P_{k}}y_{v}=\frac{1}{\omega_{y}}=\frac{1}{\omega_{z}}$
Since $\sum_{v\in V(G)}z_{v}=\sum_{v\in Supp(z)}z_{v}=1$, we get
$\sum_{v\in V_{1}}z_{v}=1-\sum_{k=2}^{\omega_{y}}\left(\sum_{v\in V_{k}}z_{v}\right)=1-(\omega_{y}-1)\frac{1}{\omega_{y}}=\frac{1}{\omega_{z}}$
Thus, for all $k\in[\omega_{z}]$,
$\sum_{v\in V_{k}}z_{v}=\frac{1}{\omega_{z}}$
Thus, $z$ satisfies 2.3, which contradicts our initial assumption. ∎
From the equality condition 1, we obtain that $y=\left\{\frac{1}{n},\frac{1}{n},\dots,\frac{1}{n}\right\}$ is a minimizer of $\Phi(G,\cdot)$. Here $Supp(y)=V(G)$. Hence, by 2.3, the graph $G$ must be complete multipartite. Let $V_{p}$ and $V_{q}$ be any two parts of $G$. Using 2.3 again, we have